calculate the temperature of a body if the resistance of a resistance thermometer is 0.5 ohms at 0 degree celsius, 55 ohms at 100degree celsius and 10 ohms at the temperature of the body
1
Expert's answer
2018-05-30T10:04:08-0400
The temperature dependence of the resistance
R(t)=R(0)(1+αt)
So
55=0.5(1+100α)
α=109/100
The resistance of a thermometer at the body temperature t
10=0.5(1+109/100 t)
19=109/100 t
Therefore
t=17.4 °C